# Trial and Trust: Addressing Byzantine Attacks with Comprehensive Defense Strategy

### v1

@@ Appendix G Local methods
$\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\left\|\nabla f_{i}(x_{i}^{t})-\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\nabla f_{i}(x_{i}^{t})\right\|^{2}$ $=$ $\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\left\|\nabla f_{i}(x_{i}^{t})-\nabla f(\overline{x}^{t})\right\|^{2}+\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\left\|\nabla f(\overline{x}^{t})-\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\nabla f_{i}(x_{i}^{t})\right\|^{2}$
$+\frac{2}{G}\sum\limits_{i\in\mathcal{G}}\left\langle\nabla f_{i}(x_{i}^{t})-\nabla f(\overline{x}^{t}),\nabla f(\overline{x}^{t})-\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\nabla f_{i}(x_{i}^{t})\right\rangle$
$=$ $\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\left\|\nabla f_{i}(x_{i}^{t})-\nabla f(\overline{x}^{t})\right\|^{2}+\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\left\|\nabla f(\overline{x}^{t})-\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\nabla f_{i}(x_{i}^{t})\right\|^{2}$
$-\frac{2}{G}\sum\limits_{i\in\mathcal{G}}\left\|\nabla f(\overline{x}^{t})-\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\nabla f_{i}(x_{i}^{t})\right\|^{2}$
@@
(${\leqslant}$ $\frac{2}{G}\sum\limits_{i\in\mathcal{G}}\left\|\nabla f_{i}(x_{i}^{t})-\nabla f_{i}(\overline{x}^{t})\right\|^{2}+\frac{2}{G}\sum\limits_{i\in\mathcal{G}}\left\|\nabla f_{i}(\overline{x}^{t})-\nabla f(\overline{x}^{t})\right\|^{2}$
$\overset{(i)}{\leqslant}$ $\frac{2}{G}\sum\limits_{i\in\mathcal{G}}\left\|\nabla f_{i}(x_{i}^{t})-\nabla f_{i}(\overline{x}^{t})\right\|^{2}+2(\delta_{1}+\delta_{2}\left\|\nabla f(\overline{x}^{t})\right\|^{2})$
(${\leqslant}$ $\frac{2}{G}\sum\limits_{i\in\mathcal{G}}\left(2L\left(f_{i}(\overline{x}^{t})-f_{i}(x_{i}^{t})-\left\langle\nabla f_{i}(x_{i}^{t}),\overline{x}^{t}-x_{i}^{t}\right\rangle\right)\right)$
$+2(\delta_{1}+\delta_{2}\|\nabla f(\overline{x}^{t})\|^{2}),$
@@
$\mathbb{E}V^{t+1}$$\leqslant$ $\mathbb{E}V^{t}-\frac{2\gamma}{G}\sum\limits_{i\in\mathcal{G}}\left\langle x_{i}^{t}-\overline{x}^{t},\nabla f_{i}(x_{i}^{t})\right\rangle+\frac{4L\gamma^{2}}{G}\sum\limits_{i\in\mathcal{G}}\left\langle x_{i}^{t}-\overline{x}^{t},\nabla f_{i}(x_{i}^{t})\right\rangle+\frac{4L\gamma^{2}}{G}\sum\limits_{i\in\mathcal{G}}\left(f_{i}(\overline{x}^{t})-f_{i}(x_{i}^{t})\right)$
@@
$\mathbb{E}V^{t+1}$$\leqslant$ $\mathbb{E}V^{t}+\frac{\gamma}{G}\sum\limits_{i\in\mathcal{G}}\left(f_{i}(\overline{x}^{t})-f_{i}(x_{i}^{t})\right)+\frac{4L\gamma^{2}}{G}\sum\limits_{i\in\mathcal{G}}\left(f_{i}(\overline{x}^{t})-f_{i}(x_{i}^{t})\right)+2\gamma^{2}(\delta_{1}+\delta_{2}\|\nabla f(\overline{x}^{t})\|^{2})+\gamma^{2}\sigma^{2}$ (9)
@@
(${\leqslant}$ $\mathbb{E}V^{t}+\frac{\gamma(1+4L\gamma)}{G}\sum\limits_{i\in\mathcal{G}}\left(\left\langle\nabla f_{i}(x_{i}^{t}),\overline{x}^{t}-x_{i}^{t}\right\rangle+\frac{L}{2}\|\overline{x}^{t}-x_{i}^{t}\|^{2}\right)+2\gamma^{2}(\delta_{1}+\delta_{2}\|\nabla f(\overline{x}^{t})\|^{2})+\gamma^{2}\sigma^{2}$
@@
$\leqslant$ $\mathbb{E}V^{t}+\frac{\gamma(1+4L\gamma)}{G}\sum\limits_{i\in\mathcal{G}}\left(\left\langle\nabla f_{i}(\overline{x}^{t})-\nabla f(\overline{x}^{t}),\overline{x}^{t}-x_{i}^{t}\right\rangle+\left\langle\nabla f(\overline{x}^{t}),\overline{x}^{t}-x_{i}^{t}\right\rangle+\frac{L}{2}\|\overline{x}^{t}-x_{i}^{t}\|^{2}\right)$
$\overset{(i)}{\leqslant}$ $\mathbb{E}V^{t}+\frac{\gamma(1+4L\gamma)}{G}\sum\limits_{i\in\mathcal{G}}\left(\frac{1}{2L}\|\nabla f_{i}(\overline{x}^{t})-\nabla f(\overline{x}^{t})\|^{2}+L\|\overline{x}^{t}-x_{i}^{t}\|^{2}+\frac{L}{2}\|\overline{x}^{t}-x_{i}^{t}\|^{2}+f(\overline{x}^{t})-f(x_{i}^{t})\right)$
$+2\gamma^{2}(\delta_{1}+\delta_{2}\|\nabla f(\overline{x}^{t})\|^{2})+\gamma^{2}\sigma^{2},$
@@
$\leqslant$ $\delta_{2}\gamma\left(2\gamma+\frac{1}{L}\right)(1+3L\gamma)^{l-1}\sum\limits_{j=t_{k\cdot l}}^{t-1}\mathbb{E}\left\|\nabla f(\overline{x}^{j})\right\|^{2}+\left(\delta_{1}\gamma\left(2\gamma+\frac{1}{L}\right)+\gamma^{2}\sigma^{2}\right)(l-1)(1+3L\gamma)^{l-1}.$
Now we tune $\gamma\leqslant\frac{1}{4(l-1)L}$. Note it is smallest of all previous $\gamma$, since $l\geqslant 2$ and consequently all previous transitions hold true. In that, way, using $\left(1+3L\gamma\right)^{l-1}\leqslant\left(1+\frac{3}{4(l-1)}\right)^{l-1}\leqslant\left(1+\frac{1}{l-1}\right)^{l-1}\leqslant 3$,
$\mathbb{E}V^{t}$$\leqslant$ $3\delta_{2}\gamma\left(\frac{1}{2(l-1)L}+\frac{1}{L}\right)\sum\limits_{j=t_{k\cdot l}}^{t-1}\mathbb{E}\left\|\nabla f(\overline{x}^{j})\right\|^{2}+3\left(\delta_{1}\gamma\left(\frac{1}{2(l-1)L}+\frac{1}{L}\right)+\gamma^{2}\sigma^{2}\right)(l-1)$
@@
Under Assumptions 2.1, 2.2(c), 2.3, 2.4, 2.5, at each iteration $t$ of Algorithm 7, the following estimation is valid:
$\mathbb{E}\hat{f}\left(\overline{x}^{t+1}\right)$$\leqslant$ $\mathbb{E}\hat{f}\left(\overline{x}^{t}\right)-\frac{\gamma}{2}\left(\left(1-4\delta_{2}\right)-3L\gamma(1+2\delta_{2})\right)\mathbb{E}\left\|\nabla f\left(\overline{x}^{t}\right)\right\|^{2}+L^{2}\gamma(1+3L\gamma)\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\mathbb{E}\left\|x_{i}^{t}-\overline{x}^{t}\right\|^{2}$
$+\frac{3L\gamma^{2}}{2G}\sigma^{2}+\gamma(2+3L\gamma)\delta_{1}+\gamma\zeta(N)+\delta.$
@@
$\hat{f}(x^{t+1})~{}$ $~{}~{}\leqslant~{}\min_{\omega\in\Delta_{1}^{n}}\hat{f}\left[\sum_{i=1}^{n}\omega_{i}\left(x^{t}-\gamma g_{i}^{t}\right)\right]+\delta.$
@@
$\hat{f}\left(\overline{x}^{t+1}\right)$ $=$ $\hat{f}(x^{t+1})\leqslant~{}\min_{\omega\in\Delta_{1}^{n}}\hat{f}\left[\sum_{i=1}^{n}\omega_{i}\left(x_{i}^{t}-\gamma g_{i}^{t}\right)\right]+\delta$
$\leqslant$ $\hat{f}\left[\frac{1}{G}\sum\limits_{i\in\mathcal{G}}x_{i}^{t}-\frac{\gamma}{G}\sum_{i\in\mathcal{G}}g_{i}^{t}\right]+\delta$
(${\leqslant}$ $\hat{f}\left(\overline{x}^{t}\right)-\left\langle\nabla\hat{f}\left(\overline{x}^{t}\right),\frac{\gamma}{G}\sum_{i\in\mathcal{G}}g_{i}^{t}\right\rangle+\frac{L\gamma^{2}}{2}\left\|\frac{1}{G}\sum_{i\in\mathcal{G}}g_{i}^{t}\right\|^{2}+\delta.$
@@
$\mathbb{E}\hat{f}\left(\overline{x}^{t+1}\right)$$\leqslant$ $\mathbb{E}\hat{f}\left(\overline{x}^{t}\right)-\left\langle\nabla\hat{f}\left(\overline{x}^{t}\right),\frac{\gamma}{G}\sum_{i\in\mathcal{G}}\nabla f_{i}(x_{i}^{t})\right\rangle+\frac{L\gamma^{2}}{2}\mathbb{E}\left\|\frac{1}{G}\sum_{i\in\mathcal{G}}g_{i}^{t}\right\|^{2}+\delta$
@@
$-\left\langle\nabla f\left(\overline{x}^{t}\right)-\nabla\hat{f}\left(\overline{x}^{t}\right),\frac{\gamma}{G}\sum_{i\in\mathcal{G}}\nabla f_{i}(x_{i}^{t})\right\rangle+\frac{L\gamma^{2}}{2}\mathbb{E}\left\|\frac{1}{G}\sum_{i\in\mathcal{G}}g_{i}^{t}\right\|^{2}+\delta$
@@
$+\frac{\gamma}{2}\left\|\frac{1}{G}\sum_{i\in\mathcal{G}}\nabla f_{i}(x_{i}^{t})\right\|^{2}+\frac{\gamma}{2}\left\|\nabla\hat{f}\left(\overline{x}^{t}\right)-\nabla f\left(\overline{x}^{t}\right)\right\|^{2}+\frac{L\gamma^{2}}{2}\mathbb{E}\left\|\frac{1}{G}\sum_{i\in\mathcal{G}}g_{i}^{t}\right\|^{2}+\delta$
(${\leqslant}$ $\mathbb{E}\hat{f}\left(\overline{x}^{t}\right)-\frac{\gamma}{2}\left\|\nabla f\left(\overline{x}^{t}\right)\right\|^{2}+\frac{\gamma}{2}\left\|\nabla f\left(\overline{x}^{t}\right)-\frac{1}{G}\sum_{i\in\mathcal{G}}\nabla f_{i}(x_{i}^{t})\right\|^{2}$
$+\gamma\left\|\nabla\hat{f}\left(\overline{x}^{t}\right)-\nabla f_{1}\left(\overline{x}^{t}\right)\right\|^{2}+\gamma\left\|\nabla f_{1}\left(\overline{x}^{t}\right)-\nabla f\left(\overline{x}^{t}\right)\right\|^{2}+\frac{L\gamma^{2}}{2}\mathbb{E}\left\|\frac{1}{G}\sum_{i\in\mathcal{G}}g_{i}^{t}\right\|^{2}+\delta$
@@
$+\gamma\left\|\nabla f_{1}\left(\overline{x}^{t}\right)-\nabla f\left(\overline{x}^{t}\right)\right\|^{2}+\frac{L\gamma^{2}}{2}\mathbb{E}\left\|\frac{1}{G}\sum_{i\in\mathcal{G}}g_{i}^{t}\right\|^{2}+\gamma\zeta(N)+\delta$
(${\leqslant}$ $\mathbb{E}\hat{f}\left(\overline{x}^{t}\right)-\frac{\gamma}{2}\left\|\nabla f\left(\overline{x}^{t}\right)\right\|^{2}+\frac{\gamma}{2}\left\|\nabla f\left(\overline{x}^{t}\right)-\frac{1}{G}\sum_{i\in\mathcal{G}}\nabla f_{i}(x_{i}^{t})\right\|^{2}$
@@
$+\frac{3L\gamma^{2}}{2}\mathbb{E}\left\|\frac{1}{G}\sum_{i\in\mathcal{G}}\nabla f_{i}(x_{i}^{t})-\nabla f(\overline{x}^{t})\right\|^{2}+\frac{3L\gamma^{2}}{2}\mathbb{E}\left\|\nabla f(\overline{x}^{t})\right\|^{2}+\gamma\zeta(N)+\delta.$
@@
$\mathbb{E}\hat{f}\left(\overline{x}^{t+1}\right)$$\leqslant$ $\mathbb{E}\hat{f}\left(\overline{x}^{t}\right)-\frac{\gamma}{2}\left(1-3L\gamma\right)\mathbb{E}\left\|\nabla f\left(\overline{x}^{t}\right)\right\|^{2}+\frac{\gamma}{2}(1+3L\gamma)\mathbb{E}\left\|\frac{1}{G}\sum_{i\in\mathcal{G}}\nabla f_{i}(x_{i}^{t})-\nabla f(\overline{x}^{t})\right\|^{2}$
$+\gamma\mathbb{E}\left\|\nabla f_{1}\left(\overline{x}^{t}\right)-\nabla f\left(\overline{x}^{t}\right)\right\|^{2}+\frac{3L\gamma^{2}}{2G}\sigma^{2}+\gamma\zeta(N)+\delta$
(${\leqslant}$ $\mathbb{E}\hat{f}\left(\overline{x}^{t}\right)-\frac{\gamma}{2}\left(1-3L\gamma\right)\mathbb{E}\left\|\nabla f\left(\overline{x}^{t}\right)\right\|^{2}+\gamma(1+3L\gamma)\mathbb{E}\left\|\frac{1}{G}\sum_{i\in\mathcal{G}}\nabla f_{i}(x_{i}^{t})-\frac{1}{G}\sum_{i\in\mathcal{G}}\nabla f_{i}(\overline{x}^{t})\right\|^{2}$
$+\gamma(1+3L\gamma)\mathbb{E}\left\|\frac{1}{G}\sum_{i\in\mathcal{G}}\nabla f_{i}(\overline{x}^{t})-\nabla f(\overline{x}^{t})\right\|^{2}+\gamma\mathbb{E}\left\|\nabla f_{1}\left(\overline{x}^{t}\right)-\nabla f\left(\overline{x}^{t}\right)\right\|^{2}+\frac{3L\gamma^{2}}{2G}\sigma^{2}$
$+\gamma\zeta(N)+\delta$
(${\leqslant}$ $\mathbb{E}\hat{f}\left(\overline{x}^{t}\right)-\frac{\gamma}{2}\left(1-3L\gamma\right)\mathbb{E}\left\|\nabla f\left(\overline{x}^{t}\right)\right\|^{2}+\frac{\gamma(1+3L\gamma)}{G}\sum_{i\in\mathcal{G}}\mathbb{E}\left\|\nabla f_{i}(x_{i}^{t})-\nabla f_{i}(\overline{x}^{t})\right\|^{2}$
$+\frac{\gamma(1+3L\gamma)}{G}\sum_{i\in\mathcal{G}}\mathbb{E}\left\|\nabla f_{i}(\overline{x}^{t})-\nabla f(\overline{x}^{t})\right\|^{2}+\gamma\mathbb{E}\left\|\nabla f_{1}\left(\overline{x}^{t}\right)-\nabla f\left(\overline{x}^{t}\right)\right\|^{2}+\frac{3L\gamma^{2}}{2G}\sigma^{2}$
$+\gamma\zeta(N)+\delta.$
@@
$+\quad\frac{\gamma(1+3L\gamma)}{G}\sum_{i\in\mathcal{G}}\mathbb{E}\left\|\nabla f_{i}(x_{i}^{t})-\nabla f_{i}(\overline{x}^{t})\right\|^{2}+\frac{3L\gamma^{2}}{2G}\sigma^{2}+\gamma(2+3L\gamma)\delta_{1}+\gamma\zeta(N)+\delta$
(${\leqslant}$ $\mathbb{E}\hat{f}\left(\overline{x}^{t}\right)-\frac{\gamma}{2}\left(\left(1-4\delta_{2}\right)-3L\gamma(1+2\delta_{2})\right)\mathbb{E}\left\|\nabla f\left(\overline{x}^{t}\right)\right\|^{2}+L^{2}\gamma(1+3L\gamma)\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\mathbb{E}\left\|x_{i}^{t}-\overline{x}^{t}\right\|^{2}$
$+\frac{3L\gamma^{2}}{2G}\sigma^{2}+\gamma(2+3L\gamma)\delta_{1}+\gamma\zeta(N)+\delta.$
@@
$\hat{f}(\overline{x}^{t+1})$ $=$ $\hat{f}\left[\frac{1}{G}\sum\limits_{i=1}^{n}x_{i}^{t}-\frac{\gamma}{G}\sum\limits_{i=1}^{n}g_{i}^{t}\right]\leqslant\hat{f}\left[\frac{1}{G}\sum\limits_{i=1}^{n}x_{i}^{t}-\frac{\gamma}{G}\sum\limits_{i=1}^{n}g_{i}^{t}\right]+\delta.$
@@
Under Assumptions 2.1, 2.2(b), 2.3, 2.4 with $\delta_{2}\leqslant\frac{1}{8}$, 2.5, for solving the problem (1), for Algorithm 7 with $\gamma\leqslant\frac{1}{25(l-1)L}$, the following holds:
@@
$+13\delta_{1}+30L\gamma(l-1)\delta_{1}+5\zeta(N)+5\frac{\delta}{\gamma}.$

### v5

@@ Appendix G Local methods
$~\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\left\|\nabla f_{i}(x_{i}^{t})-\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\nabla f_{i}(x_{i}^{t})\right\|^{2}$
$~=~\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\left\|\nabla f_{i}(x_{i}^{t})-\nabla f(\overline{x}^{t})\right\|^{2}+\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\left\|\nabla f(\overline{x}^{t})-\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\nabla f_{i}(x_{i}^{t})\right\|^{2}$
$\quad~~+\frac{2}{G}\sum\limits_{i\in\mathcal{G}}\left\langle\nabla f_{i}(x_{i}^{t})-\nabla f(\overline{x}^{t}),\nabla f(\overline{x}^{t})-\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\nabla f_{i}(x_{i}^{t})\right\rangle$
$~=~\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\left\|\nabla f_{i}(x_{i}^{t})-\nabla f(\overline{x}^{t})\right\|^{2}+\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\left\|\nabla f(\overline{x}^{t})-\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\nabla f_{i}(x_{i}^{t})\right\|^{2}$
$\quad~~-\frac{2}{G}\sum\limits_{i\in\mathcal{G}}\left\|\nabla f(\overline{x}^{t})-\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\nabla f_{i}(x_{i}^{t})\right\|^{2}$
$~\leqslant~\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\left\|\nabla f_{i}(x_{i}^{t})-\nabla f(\overline{x}^{t})\right\|^{2}$
@@
$~\overset{(i)}{\leqslant}~\frac{2}{G}\sum\limits_{i\in\mathcal{G}}\left\|\nabla f_{i}(x_{i}^{t})-\nabla f_{i}(\overline{x}^{t})\right\|^{2}+2(\delta_{1}+\delta_{2}\left\|\nabla f(\overline{x}^{t})\right\|^{2})$
[…]
$\quad~~+2(\delta_{1}+\delta_{2}\|\nabla f(\overline{x}^{t})\|^{2}),$
@@
$\mathbb{E}V^{t+1}$ $\leqslant$ $\mathbb{E}V^{t}-\frac{2\gamma}{G}\sum\limits_{i\in\mathcal{G}}\left\langle x_{i}^{t}-\overline{x}^{t},\nabla f_{i}(x_{i}^{t})\right\rangle+\frac{4L\gamma^{2}}{G}\sum\limits_{i\in\mathcal{G}}\left\langle x_{i}^{t}-\overline{x}^{t},\nabla f_{i}(x_{i}^{t})\right\rangle$
$+\frac{4L\gamma^{2}}{G}\sum\limits_{i\in\mathcal{G}}\left(f_{i}(\overline{x}^{t})-f_{i}(x_{i}^{t})\right)+2\gamma^{2}(\delta_{1}+\delta_{2}\|\nabla f(\overline{x}^{t})\|^{2})+\gamma^{2}\sigma^{2}$
@@
$\mathbb{E}V^{t+1}$ $\leqslant$ $\mathbb{E}V^{t}+\frac{\gamma}{G}\sum\limits_{i\in\mathcal{G}}\left(f_{i}(\overline{x}^{t})-f_{i}(x_{i}^{t})\right)+\frac{4L\gamma^{2}}{G}\sum\limits_{i\in\mathcal{G}}\left(f_{i}(\overline{x}^{t})-f_{i}(x_{i}^{t})\right)$ (9)
@@
[…]
$\left.+\frac{L}{2}\|\overline{x}^{t}-x_{i}^{t}\|^{2}\right)+2\gamma^{2}(\delta_{1}+\delta_{2}\|\nabla f(\overline{x}^{t})\|^{2})+\gamma^{2}\sigma^{2}$
$\leqslant$ $\mathbb{E}V^{t}+\frac{\gamma(1+4L\gamma)}{G}\sum\limits_{i\in\mathcal{G}}\left(\left\langle\nabla f_{i}(\overline{x}^{t})-\nabla f(\overline{x}^{t}),\overline{x}^{t}-x_{i}^{t}\right\rangle+\left\langle\nabla f(\overline{x}^{t}),\overline{x}^{t}-x_{i}^{t}\right\rangle\right$.
$\left.+\frac{L}{2}\|\overline{x}^{t}-x_{i}^{t}\|^{2}\right)+2\gamma^{2}(\delta_{1}+\delta_{2}\|\nabla f(\overline{x}^{t})\|^{2})+\gamma^{2}\sigma^{2}$
$\overset{(i)}{\leqslant}$ $\mathbb{E}V^{t}+\frac{\gamma(1+4L\gamma)}{G}\sum\limits_{i\in\mathcal{G}}\left(\frac{1}{2L}\|\nabla f_{i}(\overline{x}^{t})-\nabla f(\overline{x}^{t})\|^{2}+L\|\overline{x}^{t}-x_{i}^{t}\|^{2}\right$.
$\left.+\frac{L}{2}\|\overline{x}^{t}-x_{i}^{t}\|^{2}+f(\overline{x}^{t})-f(x_{i}^{t})\right)+2\gamma^{2}(\delta_{1}+\delta_{2}\|\nabla f(\overline{x}^{t})\|^{2})+\gamma^{2}\sigma^{2},$
@@
$\leqslant$ $\delta_{2}\gamma\left(2\gamma+\frac{1}{L}\right)(1+3L\gamma)^{l-1}\sum\limits_{j=t_{k\cdot l}}^{t-1}\mathbb{E}\left\|\nabla f(\overline{x}^{j})\right\|^{2}$
$+\left(\delta_{1}\gamma\left(2\gamma+\frac{1}{L}\right)+\gamma^{2}\sigma^{2}\right)(l-1)(1+3L\gamma)^{l-1}.$
Now we tune $\gamma\leqslant\frac{1}{4(l-1)L}$. Note it is smallest of all previous $\gamma$, since $l\geqslant 2$ and consequently all previous transitions hold true. In that way, using $\left(1+3L\gamma\right)^{l-1}\leqslant\left(1+\frac{3}{4(l-1)}\right)^{l-1}\leqslant\left(1+\frac{1}{l-1}\right)^{l-1}\leqslant 3$,
$\mathbb{E}V^{t}$ $\leqslant$ $3\delta_{2}\gamma\left(\frac{1}{2(l-1)L}+\frac{1}{L}\right)\sum\limits_{j=t_{k\cdot l}}^{t-1}\mathbb{E}\left\|\nabla f(\overline{x}^{j})\right\|^{2}$
$+3\left(\delta_{1}\gamma\left(\frac{1}{2(l-1)L}+\frac{1}{L}\right)+\gamma^{2}\sigma^{2}\right)(l-1)$
@@
Under Assumptions 3.1, 3.2(c), 3.3, 3.4, 3.5, at each iteration $t$ of Algorithm 7, the following estimation is valid:
$+L^{2}\gamma(1+3L\gamma)\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\mathbb{E}\left\|x_{i}^{t}-\overline{x}^{t}\right\|^{2}+\frac{3L\gamma^{2}}{2G}\sigma^{2}+\gamma(2+3L\gamma)\delta_{1}+\gamma\zeta(N).$
@@
$\hat{f}(x^{t+1})$~ $~~\leqslant~\min_{\omega\in\Delta_{1}^{n}}\hat{f}\left[\sum_{i=1}^{n}\omega_{i}\left(x^{t}-\gamma g_{i}^{t}\right)\right].$
@@
$\hat{f}\left(\overline{x}^{t+1}\right)$ $=$ $\hat{f}(x^{t+1})\leqslant~\min_{\omega\in\Delta_{1}^{n}}\hat{f}\left[\sum_{i=1}^{n}\omega_{i}\left(x_{i}^{t}-\gamma g_{i}^{t}\right)\right]$
$\leqslant$ $\hat{f}\left[\frac{1}{G}\sum\limits_{i\in\mathcal{G}}x_{i}^{t}-\frac{\gamma}{G}\sum_{i\in\mathcal{G}}g_{i}^{t}\right]$
[…]
@@
$\mathbb{E}\hat{f}\left(\overline{x}^{t+1}\right)$ $\leqslant$ $\mathbb{E}\hat{f}\left(\overline{x}^{t}\right)-\left\langle\nabla\hat{f}\left(\overline{x}^{t}\right),\frac{\gamma}{G}\sum_{i\in\mathcal{G}}\nabla f_{i}(x_{i}^{t})\right\rangle+\frac{L\gamma^{2}}{2}\mathbb{E}\left\|\frac{1}{G}\sum_{i\in\mathcal{G}}g_{i}^{t}\right\|^{2}$
@@
$-\left\langle\nabla f\left(\overline{x}^{t}\right)-\nabla\hat{f}\left(\overline{x}^{t}\right),\frac{\gamma}{G}\sum_{i\in\mathcal{G}}\nabla f_{i}(x_{i}^{t})\right\rangle+\frac{L\gamma^{2}}{2}\mathbb{E}\left\|\frac{1}{G}\sum_{i\in\mathcal{G}}g_{i}^{t}\right\|^{2}$
@@
$+\frac{\gamma}{2}\left\|\nabla f\left(\overline{x}^{t}\right)-\frac{1}{G}\sum_{i\in\mathcal{G}}\nabla f_{i}(x_{i}^{t})\right\|^{2}+\frac{\gamma}{2}\left\|\frac{1}{G}\sum_{i\in\mathcal{G}}\nabla f_{i}(x_{i}^{t})\right\|^{2}$
$+\frac{\gamma}{2}\left\|\nabla\hat{f}\left(\overline{x}^{t}\right)-\nabla f\left(\overline{x}^{t}\right)\right\|^{2}+\frac{L\gamma^{2}}{2}\mathbb{E}\left\|\frac{1}{G}\sum_{i\in\mathcal{G}}g_{i}^{t}\right\|^{2}$
@@
$+\gamma\left\|\nabla\hat{f}\left(\overline{x}^{t}\right)-\nabla f_{1}\left(\overline{x}^{t}\right)\right\|^{2}+\gamma\left\|\nabla f_{1}\left(\overline{x}^{t}\right)-\nabla f\left(\overline{x}^{t}\right)\right\|^{2}$
$+\frac{L\gamma^{2}}{2}\mathbb{E}\left\|\frac{1}{G}\sum_{i\in\mathcal{G}}g_{i}^{t}\right\|^{2}$
[…]
$+\gamma\left\|\nabla f_{1}\left(\overline{x}^{t}\right)-\nabla f\left(\overline{x}^{t}\right)\right\|^{2}+\frac{L\gamma^{2}}{2}\mathbb{E}\left\|\frac{1}{G}\sum_{i\in\mathcal{G}}g_{i}^{t}\right\|^{2}+\gamma\zeta(N)$
[…]
@@
$+\frac{3L\gamma^{2}}{2}\mathbb{E}\left\|\frac{1}{G}\sum_{i\in\mathcal{G}}\nabla f_{i}(x_{i}^{t})-\nabla f(\overline{x}^{t})\right\|^{2}+\frac{3L\gamma^{2}}{2}\mathbb{E}\left\|\nabla f(\overline{x}^{t})\right\|^{2}+\gamma\zeta(N).$
@@
$\mathbb{E}\hat{f}\left(\overline{x}^{t+1}\right)$ $\leqslant$ $\mathbb{E}\hat{f}\left(\overline{x}^{t}\right)-\frac{\gamma}{2}\left(1-3L\gamma\right)\mathbb{E}\left\|\nabla f\left(\overline{x}^{t}\right)\right\|^{2}$
$+\frac{\gamma}{2}(1+3L\gamma)\mathbb{E}\left\|\frac{1}{G}\sum_{i\in\mathcal{G}}\nabla f_{i}(x_{i}^{t})-\nabla f(\overline{x}^{t})\right\|^{2}$
$+\gamma\mathbb{E}\left\|\nabla f_{1}\left(\overline{x}^{t}\right)-\nabla f\left(\overline{x}^{t}\right)\right\|^{2}+\frac{3L\gamma^{2}}{2G}\sigma^{2}+\gamma\zeta(N)$
[…]
$+\gamma(1+3L\gamma)\mathbb{E}\left\|\frac{1}{G}\sum_{i\in\mathcal{G}}\nabla f_{i}(x_{i}^{t})-\frac{1}{G}\sum_{i\in\mathcal{G}}\nabla f_{i}(\overline{x}^{t})\right\|^{2}$
$+\gamma(1+3L\gamma)\mathbb{E}\left\|\frac{1}{G}\sum_{i\in\mathcal{G}}\nabla f_{i}(\overline{x}^{t})-\nabla f(\overline{x}^{t})\right\|^{2}+\gamma\mathbb{E}\left\|\nabla f_{1}\left(\overline{x}^{t}\right)-\nabla f\left(\overline{x}^{t}\right)\right\|^{2}$
$+\frac{3L\gamma^{2}}{2G}\sigma^{2}+\gamma\zeta(N)$
[…]
$+\frac{\gamma(1+3L\gamma)}{G}\sum_{i\in\mathcal{G}}\mathbb{E}\left\|\nabla f_{i}(x_{i}^{t})-\nabla f_{i}(\overline{x}^{t})\right\|^{2}$
$+\frac{\gamma(1+3L\gamma)}{G}\sum_{i\in\mathcal{G}}\mathbb{E}\left\|\nabla f_{i}(\overline{x}^{t})-\nabla f(\overline{x}^{t})\right\|^{2}+\gamma\mathbb{E}\left\|\nabla f_{1}\left(\overline{x}^{t}\right)-\nabla f\left(\overline{x}^{t}\right)\right\|^{2}$
$+\frac{3L\gamma^{2}}{2G}\sigma^{2}+\gamma\zeta(N).$
@@
$+\frac{\gamma(1+3L\gamma)}{G}\sum_{i\in\mathcal{G}}\mathbb{E}\left\|\nabla f_{i}(x_{i}^{t})-\nabla f_{i}(\overline{x}^{t})\right\|^{2}+\frac{3L\gamma^{2}}{2G}\sigma^{2}$
$+\gamma(2+3L\gamma)\delta_{1}+\gamma\zeta(N)$
[…]
$+L^{2}\gamma(1+3L\gamma)\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\mathbb{E}\left\|x_{i}^{t}-\overline{x}^{t}\right\|^{2}$
$+\frac{3L\gamma^{2}}{2G}\sigma^{2}+\gamma(2+3L\gamma)\delta_{1}+\gamma\zeta(N).$
@@
$\hat{f}(\overline{x}^{t+1})$ $=$ $\hat{f}\left[\frac{1}{G}\sum\limits_{i=1}^{n}x_{i}^{t}-\frac{\gamma}{G}\sum\limits_{i=1}^{n}g_{i}^{t}\right]\leqslant\hat{f}\left[\frac{1}{G}\sum\limits_{i=1}^{n}x_{i}^{t}-\frac{\gamma}{G}\sum\limits_{i=1}^{n}g_{i}^{t}\right].$
@@
Under Assumptions 3.1, 3.2(b), 3.3, 3.4 with $\delta_{2}\leqslant\frac{1}{8}$, 3.5, for solving the problem (1), for Algorithm 7 with $\gamma\leqslant\frac{1}{25(l-1)L}$, the following holds:
@@
$+13\delta_{1}+30L\gamma(l-1)\delta_{1}+5\zeta(N).$